. Two ice skaters collide on the ice. A 39.6-kg skater moving South at 6.21 m/s collides with a 52.1-kg           skater moving East at 4.33 m/s. The two skaters entangle and move together across the ice. Determine       the magnitude and direction of their post-collision velocity
Answer #1

here,

mass of first skater , m1 = 39.6 kg

initial speed of 1 , u1 = - 6.21 j m/s

mass of second skater , m2 = 52.1 kg

initial speed of 2 , u2 = 4.33 i m/s

let the final speed after the collison be v

using conservation of momentum

m1 * u1 + m2 * u2 = ( m1 + m2) * v

39.6 * ( - 6.21 j) + 52.1 * 4.33 i = ( 39.6 + 52.1) * v

solving for v

v = (2.46 i - 2.68 j ) m/s

the magnitude of final speed , |v| = sqrt(2.46^2 + 2.68^2)

|v| = 3.64 m/s

the direction of their post-collison velocity , theta = arctan(2.68/2.46) = 47.5 degree South of East